# Mean Field Game of Optimal Tracking Portfolio

### v2

@@ Abstract
This paper studies the mean field game (MFG) problem arising from a large population competition in fund management, featuring a new type of relative performance via the benchmark tracking constraint. In the $n$-agent model, each agent can strategically inject capital to ensure that the total wealth outperforms the benchmark process, which is modeled as a linear combination of the population’s average wealth process and a market index process. That is, each agent is concerned about the performance of her competitors captured by the floor constraint. With a continuum of agents, we formulate the constrained MFG problem and transform it into an equivalent unconstrained MFG problem with a reflected state process. We establish the existence of the mean field equilibrium (MFE) using the partial differential equation (PDE) approach. Firstly, by applying the dual transform, the best response control of the representative agent can be characterized in analytical form in terms of a dual reflected diffusion process. As a novel contribution, we verify the consistency condition of the MFE in separated domains with the help of the duality relationship and properties of the dual process.
@@ 1 Introduction
In the present paper, we study the relaxed tracking portfolio mean field game (MFG) problem with capital injection. In particular, we consider a system of $n$ competitive agents who are concerned with their performance relative to the population’s average wealth. When the number of agents goes to infinity, we explore the corresponding MFG problem that is more tractable. Introduced independently by Huang et al. (2006) and Lasry and Lions (2007), MFGs have emerged as a significant research area in recent years, serving as a powerful tool for studying large population games. The existing literature on $n$-player and mean field portfolio games with relative performance mostly consider the utilities defined by comparing the agent’s own wealth to the population’s average wealth. To name a few, Espinosa and Touzi (2015) established the existence and uniqueness of a Nash equilibrium for an $n$-agent game under general utility functions and portfolio constraints. Lacker and Zariphopoulou (2019) discussed both $n$-player and MFG problems of optimal portfolio under competition and constant relative risk aversion (CRRA) or constant absolute risk aversion (CARA) relative performance criteria, deriving explicit solutions. Their work was extended by Lacker and Soret (2020) to include consumption, and by Fu and Zhou (2023) to models with random return rates and volatility. Applying the master equation approach, Souganidis and Zariphopoulou (2024) analyzed a MFG problem with relative performance, deriving the MFG solution and using it to approximate the $n$-player game equilibrium. Bo et al. (2024a) examined the equilibrium consumption under linear and multiplicative external habit formation. Bo et al. (2024b) considered a relative performance MFG problem with contagious jump risk.
@@
In contrast, we consider a new type of relative performance in the sense that each agent incorporates the population’s average wealth into the benchmark constraint. That is, each agent needs to strategically choose the investment strategy and capital injection to ensure that her own wealth process always dominates the benchmark process involving the population’s average wealth. The goal of each agent is to minimize the expectation of the accumulated capital injection. After some equivalent transformations, we need to solve a type of MFGs with state reflection at zero. The primary interest in reflected state processes stems from their ability to model hard constraints and boundary behaviors, which are ubiquitous in economic, financial, and physical systems. Unlike soft penalties that can be violated at a cost, reflection imposes a strict, binding constraint that the state cannot cross. Mathematically, this is represented by a reflection term that instantaneously “pushes” the state back into the admissible domain whenever it reaches the boundary. Stochastic control of reflected diffusions, stemming from running maximum processes or singular control, has garnered increasing attention in recent research. To name a few, Han et al. (2016) concerned the optimal pricing barrier control in a regulated financial market system modelled by a reflected diffusion process. Weerasinghe and Zhu (2016) formulated a Brownian inventory control problem with capacity expansion as an equivalent stochastic control problem with reflection, solvable through analysis of the corresponding Hamilton-Jacobi-Bellman (HJB) equation. Ferrari (2019) investigated a class of singular control problems for diffusions reflected at zero. However, the research on MFGs with reflected diffusion process remains relatively unexplored. Bayraktar et al. (2019) formulated a MFG problem with drift-controlled diffusion process with reflections for large symmetric queuing systems in heavy traffic with strategic servers. By using the master equation approach, Ricciardi (2023) handled the convergence of Nash equilibrium in an $n$-player differential game towards the optimal strategies in the MFG for drift control with reflection. Recently, Bo et al. (2025) explored a MFG problem with state-control joint law dependence and the reflected state process, in which the existence of MFEs is established by developing the compactification method and the connections between the MFG and the $n$-player game problems are also discussed.
@@
In this work, we solve the MFG problem with state reflections by using the analytical approach. In the first step, we fix a deterministic differentiable function that represents the average wealth, and study the optimization problem of the representative agent. By introducing an auxiliary controlled state process with reflection, we reformulate the control problem with fictitious singular control of capital injection under floor constraint as a unconstrained regular control problem. By means of the dual transform, probabilistic representation and stochastic flow analysis, the existence of classical solution to the HJB equation is obtained and the verification theorem on the best response control is proved. As shown in Theorem 3.4, we emphasize that the optimal portfolio process exhibits path dependence on both the wealth process $V=(V_{t})_{t\in[0,T]}$ and the benchmark process $Z=(Z_{t})_{t\in[0,T]}$. This inherent path dependence makes the direct analysis based on the original state process $V=(V_{t})_{t\in[0,T]}$ analytically intractable. Consequently, the auxiliary reflected state process $X=(X_{t})_{t\in[0,T]}$ becomes indispensable, as it facilitates the representation of optimal strategies in feedback form (see Remark 3.6).
@@
Then, we aim to verify the consistency condition and find an MFE in the second step based on the characterization of the value function and feedback function of the best response control derived in the previous step. It is sufficient to show the existence of the fixed point for the mapping that stems from the consistency condition. A fundamental challenge is the piecewise expression. The optimal strategy is not in a unified form but it takes different expression across different regions (capital injection vs. non-injection), and the boundaries of the regions also depend on the population’s state distribution. This complexity is caused by the interplay of singular control (capital injection), the benchmark constraint, and mean field interactions, transforming the consistency condition into a non-standard fixed point problem. In response, we introduce the reflected dual process, which helps to provide a probability representation of best response portfolio strategy and establish the existence and uniqueness of the fixed point of the dual variable. Accordingly, we return to the primal state variable by using the dual relationship and find the MFE in piecewise form by constructing the fixed point. We split the domain into two regions depending on the model parameters, the initial level of the wealth and benchmark processes and the time horizon: on the non-capital injection region, we derive the explicit expression of the optimal investment strategy and extra capital injection is unnecessary; on the capital injection region, not explicit but calculable representations of the optimal investment and capital injection strategies are provided. The flowchart in Figure 1 outlines the roadmap of our proposed approach, which utilizes the auxiliary dual process.
@@
The remainder of the paper is organized as follows. Section 2 introduces the finite-agent and MFG problems of the optimal tracking portfolio. Section 3 proposes an auxiliary state process with reflection and derive the associated HJB equation with Neumann boundary condition for the best response control problem, in which the best response portfolio process of the representative agent can be obtained in analytical form in terms of a reflected dual process. Section 4 verifies the consistency condition and establishes the existence of the MFE with the aid of the reflected dual process. Section 5 presents some quantitative properties and numerical examples of the MFE. The proofs of main results in previous sections are collected in Section 6.
@@ 2 Finite-Agent and MFG Problem
We now introduce the relaxed benchmark tracking problem, wherein each agent is allowed to inject capital. The objective is to minimize the associated costs while ensuring that her portfolio performance surpasses a given benchmark process. Specifically, at any time $t\in[0,T]$, it is assumed that agent $i$ can strategically inject capital at such that the total wealth $V_{t}^{i}+A_{t}^{i}$ satisfies the relaxed constraint with respect to the linear combination of the average wealth process $\overline{V}_{t}^{n}=\frac{1}{n}\sum_{i=1}^{n}V^{i}_{t}$ and the market index level $Z_{t}^{i}$ that agent $i$ is concerned. The interaction of $n$ agents occurs through the average wealth process $\overline{V}^{n}_{t}$ in the relaxed constraint, which depicts the relative performance concern by agent $i$ with respect to other peers. That is, after the injection of capital $A^{i}=(A_{t}^{i})_{t\in[0,T]}$, each agent requires the total fund account $V_{t}^{i}+A_{t}^{i}$ to stay above the benchmark process $\lambda^{i}\overline{V}_{t}^{n}+(1-\lambda^{i})Z_{t}^{i}$ with $\lambda^{i}\in[0,1]$ for all time $t\in[0,T]$. Herein, the market index process $Z^{i}=(Z_{t}^{i})_{t\in[0,T]}$ for $i=1,\ldots,n$ is described as the following geometric Brownian motion (GBM) that $Z_{0}^{i}=z\in\mathds{R}_{+}:=[0,\infty)$ and
@@
Mathematically speaking, for $i=1,\ldots,n$, agent $i$ now aims to minimize the cumulative capital injected at the terminal time $T$, that is, for $(\mathrm{v},z)\in\mathds{R}_{+}^{2}$,
$\begin{cases}\displaystyle{\rm w}^{i}(\mathrm{v},z):=\text{$\inf_{(\theta^{i},A^{i})\in\mathbb{U}}\mathbb{E}\left[A_{0}^{i}+\int_{0}^{T}e^{-\rho t}dA^{i}_{t}\Big|V^{i}_{0}=\mathrm{v},Z_{0}^{i}=z\right]$}\\[10.00002pt] \displaystyle\text{subject to\ \ $\lambda^{i}\overline{V}_{t}^{n}+(1-\lambda^{i})Z_{t}^{i}\leq V^{i}_{t}+A^{i}_{t}$, $\forall t\in[0,T]$,}\end{cases}$ (2.4)
where $\lambda^{i}\in[0,1]$ represents the competition weight of agent $i$ towards her relative performance concern and $\rho>0$ is the discount factor. The admissible control set $\mathbb{U}$ is defined as the set of processes $(\theta,A)=(\theta_{t},A_{t})_{t\in[0,T]}$ such that (i) $\theta$ is an $\mathbb{F}$-adapted real-valued process satisfying $\mathbb{E}[\int_{0}^{T}|\theta_{t}|^{2}dt]<\infty$, and (ii) $A$ is an $\mathbb{F}$-adapted, non-negative and non-decreasing process whose paths are right-continuous with left limits. Here, for technical convenience, we assume that the initial wealth of each agent and the initial level of market index process are the same constants $(\mathrm{v},z)\in\mathds{R}_{+}^{2}$.
@@
$\begin{cases}\displaystyle{\rm w}(\mathrm{v},z):=\text{$\inf_{(\theta,A)\in\mathbb{U}}\mathbb{E}\left[A_{0}+\int_{0}^{T}e^{-\rho t}dA_{t}\Big|V_{0}=\mathrm{v},Z_{0}=z\right]$}\\[10.00002pt] \displaystyle\text{subject to\ \ $\lambda\overline{V}_{t}+(1-\lambda)Z_{t}\leq V_{t}+A_{t}$, $\forall t\in[0,T]$,}\end{cases}$ (2.5)
where the wealth process $V=(V_{t})_{t\in[0,T]}$ of the representative agent satisfies that $V_{0}=\mathrm{v}\in\mathds{R}_{+}$ and
@@
For a given function $f\in{\tt{C}}([0,T])$, let $(\theta^{*,f},A^{*,f})=(\theta_{t}^{*,f},A_{t}^{*,f})_{t\in[0,T]}\in\mathbb{U}$ be the best response solution to the stochastic control problem (2.5) for the representative agent. The triplet $(f,\theta^{*,f},A^{*,f})$ is called a MFE if the strategy $(\theta^{*,f},A^{*,f})$ is the best response to (2.5) such that $\mathbb{E}[V_{t}^{*,f}]=\mathrm{v}+\int_{0}^{t}f(s)ds$ for all $t\in[0,T]$, where $V^{*,f}=(V_{t}^{*,f})_{t\in[0,T]}$ is the wealth process (2.6) under the portfolio strategy $\theta^{*,f}$.
@@
  • Step 1. For a given deterministic positive function $f\in{\tt{C}}([0,T])$, we solve the stochastic control problem (2.5) for a representative agent against the fixed environment. The best response strategy for the representative agent is denoted by $(\theta^{*,f},A^{*,f})\in\mathbb{U}$ and $V^{*,f}=(V_{t}^{*,f})_{t\in[0,T]}$ stands for the wealth process in (2.6) under the portfolio strategy $\theta^{*,f}=(\theta_{t}^{*,f})_{t\in[0,T]}$.
  • Step 2. We derive an MFE by using the consistency condition, which is equivalent to finding a fixed point $f^{*}\in{\tt{C}}([0,T])$ taking values on $\mathds{R}_{+}$ to the fixed point problem $\mathbb{E}[V_{t}^{*,f}]=\mathrm{v}+\int_{0}^{t}f(s)ds$ for all $t\in[0,T]$. Then, the expected MFE is given by $(f^{*},\theta^{*,f^{*}},A^{*,f^{*}})=(f^{*}(t),\theta_{t}^{*,f^{*}},A_{t}^{*,f^{*}})_{t\in[0,T]}$.

### v3

@@ Abstract
This paper studies the mean field game (MFG) problem arising from a large population competition in fund management, featuring a new type of relative performance via the benchmark tracking. In the $n$-player model, each agent aims to minimize the expected largest shortfall of the wealth with reference to the benchmark process, which is modeled by a linear combination of the population’s average wealth process and a market index process. With a continuum of agents, we formulate the MFG problem with a reflected state process. We establish the existence of the mean field equilibrium (MFE) using the partial differential equation (PDE) approach. Firstly, by applying the dual transform, the best response control of the representative agent can be characterized in analytical form in terms of a dual reflected diffusion process. As a novel contribution, we verify the consistency condition of the MFE in separated domains with the help of the duality relationship and properties of the dual process. Moreover, based on the MFE, we construct an approximate Nash equilibrium for the $n$-player game when the number $n$ is sufficiently large.
@@ 1 Introduction
In the present paper, we study the relaxed tracking portfolio mean field game (MFG) problem. In particular, we consider a system of $n$ competitive agents who are concerned with their performance relative to the population’s average wealth. When the number of agents goes to infinity, we explore the corresponding MFG problem that is more tractable. Introduced independently by Huang et al. (2006) and Lasry and Lions (2007), MFGs have emerged as a significant research area in recent years, serving as a powerful tool for studying large population games. The existing literature on $n$-player and mean field portfolio games with relative performance mostly consider the utilities defined by comparing the agent’s own wealth to the population’s average wealth. To name a few, Espinosa and Touzi (2015) established the existence and uniqueness of a Nash equilibrium for an $n$-agent game under general utility functions and portfolio constraints. Lacker and Zariphopoulou (2019) discussed both $n$-player and MFG problems of optimal portfolio under competition and constant relative risk aversion (CRRA) or constant absolute risk aversion (CARA) relative performance criteria, deriving explicit solutions. Their work was extended by Lacker and Soret (2020) to include consumption, and by Fu and Zhou (2023) to models with random return rates and volatility. Applying the master equation approach, Souganidis and Zariphopoulou (2024) analyzed a MFG problem with relative performance, deriving the MFG solution and using it to approximate the $n$-player game equilibrium. Bo et al. (2024a) examined the equilibrium consumption under linear and multiplicative external habit formation. Bo et al. (2024b) considered a relative performance MFG problem with contagious jump risk.
@@
In contrast, we consider a new type of relative performance using the expected largest shortfall with reference to a benchmark that incorporates the population’s average wealth and an exogenous market index process. Each agent needs to strategically choose an investment strategy to minimize this new risk measure. See also some related concepts of risk measures such as the conventional expected shortfall with respect to a terminal random variable in Pham Pham (2002) and the intra-horizon expected shortfall in Farkas et al. Farkas et al. (2021), and references therein. After some equivalent transformations, we need to solve a type of MFGs with state reflection at zero. The primary interest in reflected state processes stems from their ability to model hard constraints and boundary behaviors, which are ubiquitous in economic, financial, and physical systems. Unlike soft penalties that can be violated at a cost, reflection imposes a strict, binding constraint that the state cannot cross. Mathematically, this is represented by a reflection term that instantaneously “pushes” the state back into the admissible domain whenever it reaches the boundary. Stochastic control of reflected diffusions, stemming from running maximum processes or singular control, has garnered increasing attention in recent research. To name a few, Han et al. (2016) concerned the optimal pricing barrier control in a regulated financial market system modelled by a reflected diffusion process. Weerasinghe and Zhu (2016) formulated a Brownian inventory control problem with capacity expansion as an equivalent stochastic control problem with reflection, solvable through analysis of the corresponding Hamilton-Jacobi-Bellman (HJB) equation. Ferrari (2019) investigated a class of singular control problems for diffusions reflected at zero. However, the research on MFGs with reflected diffusion process remains relatively unexplored. Bayraktar et al. (2019) formulated a MFG problem with drift-controlled diffusion process with reflections for large symmetric queuing systems in heavy traffic with strategic servers. By using the master equation approach, Ricciardi (2023) handled the convergence of Nash equilibrium in an $n$-player differential game towards the optimal strategies in the MFG for drift control with reflection. Recently, Bo et al. (2025) explored a MFG problem with state-control joint law dependence and the reflected state process, in which the existence of MFEs is established by developing the compactification method and the connections between the MFG and the $n$-player game problems are also discussed.
@@
In this work, we solve the MFG problem with state reflections by using the analytical approach. In the first step, we fix a deterministic differentiable function that represents the average wealth, and study the optimization problem of the representative agent. We reformulate the control problem with largest shortfall by introducing an auxiliary controlled state process with reflection. By means of the dual transform, probabilistic representation and stochastic flow analysis, the existence of classical solution to the HJB equation is obtained and the verification theorem on the best response control is proved. As shown in Theorem 3.3, we emphasize that the optimal portfolio process exhibits path dependence on both the wealth process $V=(V_{t})_{t\in[0,T]}$ and the benchmark process $Z=(Z_{t})_{t\in[0,T]}$. This inherent path dependence makes the direct analysis based on the original state process $V=(V_{t})_{t\in[0,T]}$ analytically intractable. Consequently, the auxiliary reflected state process $X=(X_{t})_{t\in[0,T]}$ becomes indispensable, as it facilitates the representation of optimal strategies in feedback form (see Remark 3.5).
@@
Then, we aim to verify the consistency condition and find an MFE in the second step based on the characterization of the value function and feedback function of the best response control derived in the previous step. It is sufficient to show the existence of the fixed point for the mapping that stems from the consistency condition. A fundamental challenge is the piecewise expression. The optimal strategy is not in a unified form but it takes different expression across different regions (outperforming vs. underperforming), and the boundaries of the regions also depend on the population’s state distribution. This complexity is caused by the interplay of maximum process (largest shortfall), the benchmark tracking, and mean field interactions, transforming the consistency condition into a non-standard fixed point problem. In response, we introduce the reflected dual process, which helps to provide a probability representation of best response portfolio strategy and establish the existence and uniqueness of the fixed point of the dual variable. Accordingly, we return to the primal state variable by using the dual relationship and find the MFE in piecewise form by constructing the fixed point. We split the domain into two regions depending on the model parameters, the initial level of the wealth and benchmark processes and the time horizon: on the outperforming region, we derive the explicit expression of the optimal investment strategy; on the underperforming region, not explicit but calculable representations of the optimal investment strategy is provided. Finally, with the aid of MFE, we construct an approximate Nash equilibrium for the game when the number $n$ is sufficiently large. The flowchart in Figure 1 outlines the roadmap of our proposed approach, which utilizes the auxiliary dual process.
@@
The remainder of the paper is organized as follows. Section 2 introduces the finite-agent and MFG problems of the optimal tracking portfolio. Section 3 proposes an auxiliary state process with reflection and derive the associated HJB equation with Neumann boundary condition for the best response control problem, in which the best response portfolio process of the representative agent can be obtained in analytical form in terms of a reflected dual process. Section 4 verifies the consistency condition and establishes the existence of the MFE with the aid of the reflected dual process. Section 5 presents some quantitative properties and numerical examples of the MFE. Section 6 constructs an approximate Nash equilibrium for the game with a sufficiently large but finite number $n$ of agents. Section 7 summarizes the main results and discusses future research directions. The proofs of main results in previous sections are collected in Section 8.
@@ 2 Finite-Agent and MFG Problem
We now introduce the relaxed benchmark tracking problem based on the expected largest shortfall by allowing the wealth process to fall below the given benchmark process. The benchmark process is modeled as a linear combination of the average wealth process $\overline{V}_{t}^{n}=\frac{1}{n}\sum_{i=1}^{n}V^{i}_{t}$ and the market index level $Z_{t}^{i}$ that agent $i$ is concerned. The interaction of $n$ agents occurs through the average wealth process $\overline{V}^{n}_{t}$ which depicts the relative performance concern by agent $i$ with respect to other peers. Herein, the market index process $Z^{i}=(Z_{t}^{i})_{t\in[0,T]}$ for $i=1,\ldots,n$ is described as the following geometric Brownian motion (GBM) that $Z_{0}^{i}=z\in\mathds{R}_{+}:=[0,\infty)$ and
@@
Given the benchmark process $Z^{i}=(Z_{t}^{i})_{t\in[0,T]}$, denote by $A^{i}=(A_{t}^{i})_{t\in[0,T]}$ the largest shortfall (up to time $t$) when the wealth falls below the benchmark, that is
$A^{i}_{t}=0\vee\sup_{s\leq t}\left(\lambda^{i}\overline{V}_{s}+(1-\lambda^{i})Z_{s}^{i}-V_{s}^{i}\right),\penalty 10000\ \penalty 10000\ \forall t\in[0,T],$
where $\lambda^{i}\in[0,1]$ represents the competition weight of agent $i$ towards her relative performance concern. Then, $\mathbb{E}[A_{0}^{i}+\int_{0}^{T}e^{-\rho t}dA^{i}_{t}]$ defines a risk measure on the expected largest shortfall with reference to the benchmark. The agent needs to strategically choose the portfolio control to minimize this risk measure on tracking.
@@
Mathematically speaking, for $i=1,\ldots,n$, agent $i$ now aims to minimize the expected largest shortfall at the terminal time $T$, that is, for $(\mathrm{v},z)\in\mathds{R}_{+}^{2}$,
${\rm w}^{i}(\mathrm{v},z):=\text{$\inf_{\theta^{i}\in\mathbb{U}}\mathbb{E}\left[A_{0}^{i}+\int_{0}^{T}e^{-\rho t}dA^{i}_{t}\Big|V^{i}_{0}=\mathrm{v},Z_{0}^{i}=z\right]$}$ (2.4)
where $\rho>0$ is the discount factor. The admissible control set $\mathbb{U}$ is defined as the set of processes $\theta=(\theta_{t})_{t\in[0,T]}$ such that $\theta$ is an $\mathbb{F}$-adapted real-valued process satisfying $\mathbb{E}[\int_{0}^{T}|\theta_{t}|^{2}dt]<\infty$. Here, for technical convenience, we assume that the initial wealth of each agent and the initial level of market index process are the same constants $(\mathrm{v},z)\in\mathds{R}_{+}^{2}$.
@@
${\rm w}(\mathrm{v},z):=\inf_{\theta\in\mathbb{U}}\mathbb{E}\left[A_{0}+\int_{0}^{T}e^{-\rho t}dA_{t}\Big|V_{0}=\mathrm{v},Z_{0}=z\right].$ (2.5)
Here the wealth process $V=(V_{t})_{t\in[0,T]}$ of the representative agent satisfies that $V_{0}=\mathrm{v}\in\mathds{R}_{+}$ and
@@
The largest shortfall process $A=(A_{t})_{t\in[0,T]}$ of the representative agent is given by
$A_{t}=0\vee\sup_{s\leq t}\left(\lambda f(s)+(1-\lambda)Z_{s}-V_{s}\right),\quad\forall t\in[0,T],$
@@
For a given function $f\in{\tt{C}}([0,T])$, let $\theta^{*,f}=(\theta_{t}^{*,f})_{t\in[0,T]}\in\mathbb{U}$ be the best response solution to the stochastic control problem (2.5) for the representative agent. The triplet $(f,\theta^{*,f})$ is called a MFE if the strategy $\theta^{*,f}$ is the best response to (2.5) such that $\mathbb{E}[V_{t}^{*,f}]=\mathrm{v}+\int_{0}^{t}f(s)ds$ for all $t\in[0,T]$, where $V^{*,f}=(V_{t}^{*,f})_{t\in[0,T]}$ is the wealth process (2.6) under the portfolio strategy $\theta^{*,f}$.
@@
  • Step 1. For a given deterministic positive function $f\in{\tt{C}}([0,T])$, we solve the stochastic control problem (2.5) for a representative agent against the fixed environment. The best response strategy for the representative agent is denoted by $\theta^{*,f},\in\mathbb{U}$ and $V^{*,f}=(V_{t}^{*,f})_{t\in[0,T]}$ stands for the wealth process in (2.6) under the portfolio strategy $\theta^{*,f}=(\theta_{t}^{*,f})_{t\in[0,T]}$.
  • Step 2. We derive an MFE by using the consistency condition, which is equivalent to finding a fixed point $f^{*}\in{\tt{C}}([0,T])$ taking values on $\mathds{R}_{+}$ to the fixed point problem $\mathbb{E}[V_{t}^{*,f}]=\mathrm{v}+\int_{0}^{t}f(s)ds$ for all $t\in[0,T]$. Then, the expected MFE is given by $(f^{*},\theta^{*,f^{*}})=(f^{*}(t),\theta_{t}^{*,f^{*}})_{t\in[0,T]}$.